OpenStudy (anonymous):

I need a step by step guide for this problem: integrate e^(t^2) * t^3

myininaya (myininaya):

you could try a subsitution u=t^2

myininaya (myininaya):

recall by law of exponents you can write t^3 as t^2*t

OpenStudy (anonymous):

Correct, resulting in \[(1/2) \int\limits_{}^{}e^u * u du\]

myininaya (myininaya):

then integration by parts

OpenStudy (anonymous):

How does that go?

myininaya (myininaya):

do you know the product rule for derivatives?

OpenStudy (anonymous):

yes, (F*G)' = F'*G + F*G'

myininaya (myininaya):

\[(fg)'=f'g+fg' \\ \text{ integrate both sides } \\ fg=\int\limits_{}^{}f'g dx+\int\limits_{}^{}fg'dx \\ fg-\int\limits_{}^{}f'gdx=\int\limits_{}^{}fg' dx\]

myininaya (myininaya):

choose f to be u and choose g' to be e^u

myininaya (myininaya):

\[\int\limits_{}^{}fg' du=fg-\int\limits_{}^{}f' g du\] f=u g'=e^u f'=? g=?

myininaya (myininaya):

have you not learned integration by parts yet?

myininaya (myininaya):

if so I don't know another way to do the integral

OpenStudy (anonymous):

My physics friend was able to finish off the question. Thank you for getting me started.